# D-69120 Heidelberg, Germany

### v1

@@ Abstract
Aims. Understanding the rotational dynamics of interstellar dust grains is quintessential for the analysis of the observed dust polarization signal. We aim to constrain the set of parameters for an accurate description of the rotational spin-up process of ballistic dust grain aggregates driven by radiative torques (RATs).
@@
Results. We find that the canonical parametrization for the torque efficiency overestimates the maximal angular velocity $\omega_{\mathrm{RAT}}$ caused by RATs of grain aggregates, and to resolve this problem we propose a new parametrization which predicts $\omega_{\mathrm{RAT}}$ more accurately. We find that RATs are strongest for larger grains with smaller monomer-density, this manifests as size and monomer-density dependence in the constant part of the parametrization. Following the constant part, the parametrization has two power-laws with different slopes, which retain the universality for all grain sizes. The maximum grain rotation does not scale linearly with radiation strength, due to different drag mechanisms dominating depending on grain material and environment. The angular velocity $\omega_{\mathrm{RAT}}$ of individual single dust grains has a wide distribution and may even differ from the mean by up to two orders of magnitude.
@@
dust, microphysics, radiative torques, rotation, ballistic aggregates, RAT, porosity, alignment, spin-up
@@ 1 Introduction
Dust is a common component in astronomic environments, from galaxies to stellar discs and therefore plays a crucial role in observations. Most astronomers come into contact with dust because of dust extinction, which affects almost all observations (Draine 2003). Calzetti et al. (2000) as well as Bernstein et al. (2002) estimated, that $30\ \%$ or more of the energy produced by stars in the universe is absorbed by dust and re-emitted in the infrared. The dust helps to cool down gas, facilitating gravitational collapse (Omukai et al. 2005), also playing a major role in later stages of star formation, as well as constituting the first stage in the formation of planets (Dullemond et al. 2008; Liu & Ji 2020). Rapidly rotating dust grains are the common source for polarized light, carrying information about the magnetic field orientation (see e.g. Lazarian 2007; Andersson 2015, for review). The catalysis of molecular hydrogen (Hollenbach & Salpeter 1971; Cazaux & Tielens 2002) and other astrochemical reactions in the ISM (Dishoeck 2014) is also impacted by grain rotation (Hoang & Tram 2020). The first models trying to explain the observed dust alignment, like paramagnetic relaxation (Davis & Greenstein 1951) or the Gold mechanism (Gold 1952) both encountered problems and were ultimately unable to account for the observed alignment. A big leap in the study of alignment was made, when Purcell (1979) realized, that grains could experience systematic torques acting in a specific direction in grain coordinates for large time spans compared to the timescale it takes gas collisions to randomize the grain rotation. With such systematic torques the grain can spin up to much faster angular velocities than the typical thermal rotation rate. With suprathermal rotation, a grain can remain in a given orientation for much longer than the gas collision timescale, allowing alignment of angular momentum with the maximal moment of inertia via inelastic relaxation or the Barnett relaxation discovered by Purcell (1979), based on the Barnett. (1917) effect, which causes magnetization along the rotation axis in a paramagnetic body. If the grains rotation axis is not aligned with a principal axis of inertia, the rotation axis and therefore the magnetisation axis will precess in body coordinates, dissipating energy and aligning the rotation axis with the axis of maximal moment of inertia. In the presence of an external magnetic field the Davis-Greenstein-process can align the grain with the magnetic field (Purcell 1979).
@@
The idea of systematic radiative torques (RATs) as the cause for rapid grain rotation had already been advanced decades prior (Dolginov 1972; Dolginov & Mytrophanov 1976), but only became an integral part of dust alignment theory with the work of Draine & Weingartner (1996). Using the discrete dipole approximation (DDA) method it was shown, that scattering and absorption of light from irregular grains results in torques acting on timescales much larger than random gas collisions. Their calculations were done on a irregular shape of intersecting cuboids and it was found, that radiative torques depend strongly on grain size. Additional grain shapes were investigated in the following work (Draine & Weingartner 1997) in the context of anisotropic radiation, $\mathrm{H}_{2}$-formation torques and paramagnetic relaxation. In most cases the grain reaches long-term stable conditions by aligning the grains principal axis with the magnetic field. Later, Lazarian & Hoang (2007) (LH07 hereafter) introduced an analytic model (AMO) for a spherical grain attached to a flat rectangular mirror via a weightless non interacting rod in this way capturing the helical properties of irregular grains. The AMO was compared to DDA simulations of irregular grain models, corresponding to the grain shapes in Draine & Weingartner (1997) and found a good agreement. The alignment of grains principal axis with the direction of incoming radiation was confirmed when magnetic field effects were subdominant. The parametrization of the RAT spin-up efficiency was found to be a power-law with an exponent of about $-3$. Physical implications of RATs were explored in following studies (Lazarian & Hoang 2008; Hoang & Lazarian 2008, 2009, 2016) and the model was expanded by the addition of superparamagnetic relaxation due to iron inclusions. This increased the range of parameters in which grains can efficiently align with the magnetic field which is necessary to explain high degrees of polarization from observations (Planck Collaboration et al. 2020). The AMO was used successfully to model synthetic dust polarization observations, as long as free alignment parameters could be used to adjust the model (Bethell et al. 2007; Hoang & Lazarian 2014; Reissl et al. 2016). Herranen et al. (2019) then made a step towards testing more realistic grain shapes: gaussian ellipsoids, oblate spheroids, prolate spheroids, and spheres with random deformations and found a shallower slope of $-2.6$ for the power-law for the RAT parametrization. In this work we take one more step towards more realistic grain shapes, by simulating random typically fractal aggregate grains composed of different sized monomers and compare their RATs to previous results. The motivation is to model the grain growth ab-initio using the current understanding, that dust is originally formed in AGB stellar winds or supernovae (Todini & Ferrara 2001; Schneider et al. 2014; Nanni et al. 2013). As the dust forming materials in the winds cool down, heavy elements condense into spherical particles - in the following called monomers - which are the building blocks for larger dust grains formed by monomer accretion and coagulation of grains (Gobrecht et al. 2016; Norris et al. 2012). More realistic grain modeling may result in differing interactions with radiation. One example being the study of Arnold et al. (2019), that found significant differences in dust blowout sizes between compact and agglomerated grains.
@@
The aim of this work is to improve the understanding of RATs by simulating individual grains composed of monomers for different sources of radiation. The theoretical basis for our simulations and models for the sources are presented in Sect. 3. The properties of grains and the way they are modeled is presented in Sect. 2, usage of the simulation codes is explained in Sect. 5, results are shown in Sect. 6 and discussed in Sect. 7, a summary and outlook are given in Sect. 8.
@@ 3 Maximal grain rotation
For our grain dynamics modelling we assume stable alignment of the grains rotation with the direction of incoming radiation. For a single grain stably rotating about its principal axis the maximal possible angular momentum $\omega_{\mathrm{RAT}}$ caused by RATs can be calculated as
@@
S1 50 $1.5\cdot 10^{4}$ $4.6\cdot 10^{5}$ 0.7 $5.6\cdot 10^{2}$
S2 80 $2.5\cdot 10^{4}$ $9.0\cdot 10^{6}$ 0.9 $1.1\cdot 10^{4}$
AGN - - $10^{10}$ 1.0 $1.2\cdot 10^{7}$
Table 2: Source parameters: radius $R$ and effective temperature $T_{\mathrm{eff}}$ of stellar sources S1 and S2, luminosity $L_{\mathrm{bol}}$ of localized sources, anisotropy $\gamma$ and finally radiation energy density $u$ given in relation to $u_{\mathrm{ISRF}}=8.64\cdot 10^{-14}\ \mathrm{J\ m^{-3}}$. The distance to the localized sources S1, S2 and AGN, is taken to be $1\,\mathrm{pc}$.
@@ 4 The radiation field
Typical sources of radiation are chosen to evaluate the RATs they cause by iradiating the grains. Among the different sources are two O-type stars labeled S1 and S2, respectively, with a flux of
@@ 5 Optical and mechanical dust properties
To test the predictions of the RAT theory (Lazarian & Hoang 2007; Hoang et al. 2014) we consider silicates (SC) and carbonaceous (CA) grains, typical materials present in the ISM (see e.g. Andersson 2015, for review). The optical properties of the materials are represented by the wavelength-dependent complex refractive index ${\bar{n}=n+ik}$. With $n$ being the real part describing refraction of light interacting with the dust grain and the imaginary part $k$ governs the light attenuation. For the SC take the (Astro)silicate model data as presented in Draine & Lee (1984); Laor & Draine (1993) and the optical properties of hydrogenated amorphous carbon are outlined in Jones (2012) 222 In detail: The refractive indices of the CA material is taken from http://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/542/A98 for $30\ \mathrm{nm}$ hydrocarbon particles with $E_{g}=2.67\,\mathrm{eV}$ and the SC-data is listed as eps_Sil in https://www.astro.princeton.edu/~draine/dust/dust.diel.html. Ice mantles were not considered here, as high intensity radiation environments would not allow them. Assuming condensation of all ice species at a temperature above $150\,\mathrm{K}$ (Zhang et al. 2015) and comparing with dust temperatures in Tab. 3, we would expect ice-free grains for the AGN source and for CA grains around the stellar sources. The ice-free assumption may therefore generally not be valid for the ISRF and SC grains around the stellar sources. Additionally comparing with these dust temperatures assumes there is no significant difference in dust temperature when a icy surface is considered instead of the silicate/corbonaceous surface.
@@
where $\hat{e}_{\mathrm{1}}$ is anti-parallel to the wave-vector of incoming light. For simplicity, we assume that the grain is aligned with the direction of radiation i.e. $\theta=0^{\circ}$. Subsequently, the direction of radiation and rotation coincide ($\hat{e}_{\mathrm{1}}\ ||\ \hat{a}_{\mathrm{1}}$) and the RAT efficiency factor becomes ${Q}_{\mathrm{\Gamma}}\equiv Q_{\mathrm{e1}}\left(\theta\right)\hat{e}_{\mathrm{1}}$
Simulation results are averaged over 21 equidistant values of the rotation angle of the grain around $\hat{a}_{1}$. The results are computed for 50 wavelengths between $0.05\,\mu\mathrm{m}$ and $150\,\mu\mathrm{m}$ within equidistant bins in log-space. Each combination of material, monomer-density, size, and shape (see Sect. 2) is simulated with five different resolutions i.e. the separation of two nearest dipoles. Those are combined using an average weighted by dipole resolution. In addition to the RAT efficiency factor ${Q}_{\mathrm{\Gamma}}$, DDSCAT also provides the efficiency of absorption ${Q}_{\mathrm{abs}}$ required to evaluate the photon-drag-timescale $\tau_{\mathrm{IR}}$ and the dust temperature $T_{\mathrm{dust}}$. The gas-torque-efficiency $Q_{\mathrm{gas}}$ needed to evaluate the gas-drag-timescale $\tau_{\mathrm{gas}}$ is calculated with a Monte Carlo (MC) based gas-dust collision approach as outlined in Reissl et al. (2022). The corresponding gas parameters [^3] We note that the gas number density $n_{\mathrm{gas}}$ is lower than in a typical AGN torus. However, we can position the dust grain outside the torus, selecting a position in which $n_{\mathrm{gas}}$ is similar to the ISM, e.g. at higher inclinations from the plane of the torus. are listed in Tab. 1.
CA SC
source $T_{\mathrm{min}}$ [K] $T_{\mathrm{mean}}$ [K] $T_{\mathrm{max}}$ [K] $T_{\mathrm{min}}$ [K] $T_{\mathrm{mean}}$ [K] $T_{\mathrm{max}}$ [K]
AGN 780 925 1164 312 375 484
S2 274 302 343 81 91 106
S1 186 197 212 45 49 55
ISRF 72 79 88 15 16 18
Table 3: Range of the resulting dust temperatures $T_{\mathrm{dust}}$ considering SC and CA materials and the different source AGN, S1, S2, and ISRF, respectively.
@@ 6.1 Drag timescales
Resulting timescales for the gas-drag and the photon-drag, respectively, taking different gas densities and dust temperatures are shown in Fig. 3. The gas-drag-timescale $\tau_{\mathrm{gas}}$ is presented in the left panel as an average of SC and CA materials. We emphasize that the gas-drag-efficiency $Q_{\mathrm{gas}}$ is only dependent on grain shape in our modelling and is independent of material. We see that denser gas equals a smaller $\tau_{\mathrm{gas}}$ in agreement with the literature (see e.g. Draine & Weingartner 1997 and Reissl et al. 2023). The dependence on monomer-density is marginal, with the BA grains having a stronger gas-drag than the BAM1 and BAM2 grains. This is due to the dependence of $Q_{\mathrm{gas}}$ on the surface area.
In detail, the BA grains present a larger surface area for interaction with gas particles. Following Eq.5, the expected size dependence of $\tau_{\mathrm{gas}}$ would be linear ($I_{\mathrm{max}}\propto a_{\mathrm{eff}}^{5}$). However it is almost constant above $100\,\mathrm{nm}$, which is due to $Q_{\mathrm{gas}}$ increasing linearly with size. In principle the gas drag efficiency $Q_{\mathrm{gas}}$ should only depend on the shape of the grain and not the size. That $Q_{\mathrm{gas}}$ nonetheless increases with size may be explained by $a_{\mathrm{eff}}$ not being well suited to describe grain size in this situation. The calculation of $Q_{\mathrm{gas}}$ uses the simulated gas-drag-torque, which depends on the angular momentum lost by the grain when a gas particle leaves the surface. This in turn scales linearly with the distance from the axis of rotation of the grain from where the gas particle is ejected. For porous grains this distance is larger than $a_{\mathrm{eff}}$. The torque simulated in this way contains the information about the true shape of the grain, however, in the formula for $Q_{\mathrm{gas}}$ the size $a_{\mathrm{eff}}$ is used, which creates the size dependence of $Q_{\mathrm{gas}}$.
The photon-drag-timescale is shown in the middle for SC grains and on the right for CA grains in Fig. 3. As the efficiency of absorption $Q_{\mathrm{abs}}$ is overall larger for SC, more photons are absorbed and emitted causing a stronger photon-drag. The number of emitted photons also depends on dust temperature, with hotter dust experiencing stronger drag (see also Tab. 3). The dependence on monomer-density is marginal, and is related to the fact that the moment of inertia being larger in the case of BA grains than for the more compact BAM2 grains given the same effective radius (see Appendix A in Reissl et al. 2023).
@@
Figure 6: The same as Fig. 5 but for the average torque efficiency $Q_{\Gamma}$ over $\lambda/a_{\mathrm{eff}}$ in comparison with the conical power-law parameterizaiton (red dotted line) as suggested in LH07.
Figure 7: Two-power law parametrization shown as a red line for CA (left) and SC (right) BA-grains in comparison with the full set of the numerical DDSCAT simulations (black dots) for grains of size $a_{\mathrm{eff}}=200\ \mathrm{nm}$ where the absolute value of the torque efficiency $|Q_{\Gamma}|$ are plotted as a function of $\lambda/a_{\mathrm{eff}}$. The constant part $c$ is fitted for each grain size up to a value of $\lambda/a_{\mathrm{eff}}=3$, the steeper power-law is fitted as a linear function to the logarithm of the data for all sizes combined in the interval $103<\lambda/a_{\mathrm{eff}}<$ 10, same for the shallow power-law in the interval $\lambda/a_{\mathrm{eff}}>20$. Finally, the two intersection points $a$ and $b$ are calculated and used to create a continuous parametrization for $|Q_{\Gamma}|$ (see Eq. 13).

### v2

@@ Abstract
Aims. Understanding the rotational dynamics of interstellar dust grains is quintessential for the analysis of the observed dust polarization signal. Additionally, high rotation velocities may cause rotational disruption of the grains, impacting the grain size distribution. We aim to constrain the set of parameters for an accurate description of the rotational spin-up process of ballistic dust grain aggregates driven by radiative torques (RATs).
@@
Results. We find that the canonical parametrization for the torque efficiency overestimates the maximal angular velocity $\omega_{\mathrm{RAT}}$ caused by RATs acting on ballistic grain aggregates, and to resolve this problem we propose a new parametrization which predicts $\omega_{\mathrm{RAT}}$ more accurately. We find that RATs are most efficient for larger grains with smaller monomer-density, this manifests as size and monomer-density dependence in the constant part of the parametrization. Following the constant part, the parametrization has two power-laws with different slopes, which retain the universality for all grain sizes. The maximum grain rotation does not scale linearly with radiation strength, due to different drag mechanisms dominating depending on grain material and environment. The angular velocity $\omega_{\mathrm{RAT}}$ of individual single dust grains has a wide distribution and may even differ from the mean by up to two orders of magnitude. Even though ballistic aggregates have a lower RAT efficiency, strong sources of radiation (larger than $\approx 100$ times the typical interstellar radiation field) may still produce rotation velocities large enough to cause rotational disruption of dust grains.
@@
dust, microphysics, radiative torques, RAT, rotation, ballistic aggregates, porous grains, drag, spin-up, alignment, disruption
@@ 1 Introduction
Dust is a common component in astronomic environments, from galaxies to stellar discs and therefore plays a crucial role in observations. Most astronomers come into contact with dust because of dust extinction, which affects almost all observations (Draine 2003). Calzetti et al. (2000) as well as Bernstein et al. (2002) estimated, that $30\ \%$ or more of the energy produced by stars in the universe is absorbed by dust and re-emitted in the infrared. The dust helps to cool down gas, facilitating gravitational collapse (Omukai et al. 2005), also playing a major role in later stages of star formation, as well as constituting the first stage in the formation of planets (Dullemond et al. 2008; Liu & Ji 2020). Rapidly rotating dust grains are the common source for polarized light, carrying information about the magnetic field orientation (see e.g. Lazarian 2007; Andersson et al. 2015, for review). The catalysis of molecular hydrogen (Hollenbach & Salpeter 1971; Cazaux & Tielens 2002) and other astrochemical reactions in the ISM (Dishoeck 2014) is also impacted by grain rotation (Hoang & Tram 2020). The first models trying to explain the observed dust alignment, like paramagnetic relaxation (Davis & Greenstein 1951) or the Gold mechanism (Gold 1952) both encountered problems and were ultimately unable to account for the observed alignment. A big leap in the study of alignment was made, when Purcell (1979) realized, that grains could experience systematic torques acting in a specific direction in grain coordinates for large time spans compared to the timescale it takes gas collisions to randomize the grain rotation. With such systematic torques the grain can spin up to much faster angular velocities than the typical thermal rotation rate. With suprathermal rotation, a grain can remain in a given orientation for much longer than the gas collision timescale, allowing alignment of angular momentum with the maximal moment of inertia via inelastic relaxation or the Barnett relaxation discovered by Purcell (1979), based on the Barnett (1917) effect, which causes magnetization along the rotation axis in a paramagnetic body. If the grains rotation axis is not aligned with a principal axis of inertia, the rotation axis and therefore the magnetization axis will precess in body coordinates, dissipating energy and aligning the rotation axis with the axis of maximal moment of inertia. In the presence of an external magnetic field the Davis-Greenstein-process can align the grain with the magnetic field (Purcell 1979).
@@
The idea of systematic radiative torques (RATs) as the cause for rapid grain rotation had already been advanced decades prior (Dolginov 1972; Dolginov & Mytrophanov 1976), but only became an integral part of dust alignment theory with the work of Draine & Weingartner (1996). Using the discrete dipole approximation (DDA) method it was shown, that scattering and absorption of light from irregular grains results in torques acting on timescales much larger than random gas collisions. Their calculations were done on a irregular shape of intersecting cuboids and it was found, that radiative torques depend strongly on grain size. Additional grain shapes were investigated in the following work (Draine & Weingartner 1997) in the context of anisotropic radiation, $\mathrm{H}_{2}$-formation torques and paramagnetic relaxation. In most cases the grain reaches long-term stable conditions by aligning the grains principal axis with the magnetic field. Later, Lazarian & Hoang (2007a) (LH07 hereafter) introduced an analytic model (AMO) for a spherical grain attached to a flat rectangular mirror via a weightless non interacting rod in this way capturing the helical properties of irregular grains. The AMO was compared to DDA simulations of irregular grain models, corresponding to the grain shapes in Draine & Weingartner (1997) and found a good agreement. The alignment of grains principal axis with the direction of incoming radiation was confirmed when magnetic field effects were subdominant. The parametrization of the RAT spin-up efficiency was found to be a power-law with an exponent of about $-3$. Physical implications of RATs were explored in following studies (Lazarian & Hoang 2008; Hoang & Lazarian 2008, 2009, 2016) and the model was expanded by the addition of superparamagnetic relaxation due to iron inclusions. This increased the range of parameters in which grains can efficiently align with the magnetic field which is necessary to explain high degrees of polarization from observations (Planck Collaboration et al. 2020). The AMO was used successfully to model synthetic dust polarization observations, as long as free alignment parameters could be used to adjust the model (Bethell et al. 2007; Hoang & Lazarian 2014; Reissl et al. 2016). Herranen et al. (2019) then made a step towards testing more realistic grain shapes: gaussian ellipsoids, oblate spheroids, prolate spheroids, and spheres with random deformations and found a shallower slope of $-2.6$ for the power-law for the RAT parametrization. In this work we take one more step towards more realistic grain shapes, by simulating random typically fractal aggregate grains composed of different sized monomers and compare their RATs to previous results. The motivation is to model the grain growth ab-initio using the current understanding, that dust is originally formed in AGB stellar winds or supernovae (Todini & Ferrara 2001; Schneider et al. 2014; Nanni et al. 2013). As the dust forming materials in the winds cool down, heavy elements condense into spherical particles - in the following called monomers - which are the building blocks for larger dust grains formed by monomer accretion and coagulation of grains (Gobrecht et al. 2016; Norris et al. 2012). More realistic grain modeling may result in differing interactions with radiation. One example being the study of Arnold et al. (2019), that found significant differences in dust blowout sizes between compact and agglomerated grains.
@@
The aim of this work is to improve the understanding of RATs by simulating individual grains composed of monomers for different sources of radiation. The properties of grains and the way they are modeled is presented in Sect. 2. The theoretical basis for our simulations are presented in Sect. 3 and models for the sources in Sect. 4, usage of the simulation codes is explained in Sect. 5, results are shown in Sect. 6 and discussed in Sect. 7, a summary and outlook are given in Sect. 8.
@@ 3 Maximal grain rotation
For our grain dynamics modeling we assume stable alignment of the grains rotation with the direction of incoming radiation. For a single grain stably rotating about its principal axis the maximal possible angular momentum $\omega_{\mathrm{RAT}}$ caused by RATs can be calculated as
@@
S1 50 $1.5\cdot 10^{4}$ $4.6\cdot 10^{5}$ 0.7 $5.6$
S2 80 $2.5\cdot 10^{4}$ $9.0\cdot 10^{6}$ 0.9 $1.1\cdot 10^{2}$
AGN - - $10^{10}$ 1.0 $1.2\cdot 10^{5}$
 [^2] Includes: stellar radius $R$ and effective temperature $T_{\mathrm{eff}}$ for the stellar sources S1 and S2, bolometric luminosity $L_{\mathrm{bol}}$ for all localized sources, anisotropy $\gamma$ and radiation energy density $u$ given in relation to $u_{\mathrm{ISRF}}=8.64\cdot 10^{-14}\ \mathrm{J\ m^{-3}}$. The distance to the localized sources S1, S2 and AGN, is taken to be $10\,\mathrm{pc}$.
@@ 4 The radiation field
Typical sources of radiation are chosen to evaluate the RATs they cause by irradiating the grains. Among the different sources are two O-type stars labeled S1 and S2, respectively, with a flux of
@@ 5 Optical and mechanical dust properties
To test the predictions of the RAT theory (LH07; Hoang et al. 2014) we consider silicates (SC) and carbonaceous (CA) grains, typical materials present in the ISM (see e.g. Draine 2003, for review). The optical properties of the materials are represented by the wavelength-dependent complex refractive index ${\bar{n}=n+ik}$. With $n$ being the real part describing refraction of light interacting with the dust grain and the imaginary part $k$ governs the light attenuation. For the SC take the (Astro)silicate model data as presented in Draine & Lee (1984); Laor & Draine (1993) and the optical properties of hydrogenated amorphous carbon are outlined in Jones (2012) 333 In detail: The refractive indices of the CA material is taken from http://vizier.cds.unistra.fr/viz-bin/VizieR?-source=J/A+A/542/A98 for $30\ \mathrm{nm}$ hydrocarbon particles with $E_{g}=2.67\,\mathrm{eV}$ and the SC-data is listed as eps_Sil in https://www.astro.princeton.edu/~draine/dust/dust.diel.html. Ice mantles were not considered here, as high intensity radiation environments would not allow them. Assuming condensation of all ice species at a temperature above $150\,\mathrm{K}$ (Zhang et al. 2015) and comparing with dust temperatures in Tab. 3, we would expect ice-free grains for the AGN source and for CA grains around the stellar sources. The ice-free assumption may therefore generally not be valid for the ISRF and SC grains around the stellar sources. Additionally comparing with these dust temperatures assumes there is no significant difference in dust temperature when a icy surface is considered instead of the silicate/carbonaceous surface.
@@
where $\hat{e}_{\mathrm{1}}$ is anti-parallel to the wave-vector of incoming light. For simplicity, we assume that the grain is aligned with the direction of radiation i.e. $\Theta=0^{\circ}$. Subsequently, the direction of radiation and rotation coincide ($\hat{e}_{\mathrm{1}}\ ||\ \hat{a}_{\mathrm{1}}$) and the RAT efficiency factor becomes ${Q}_{\mathrm{\Gamma}}\equiv Q_{\mathrm{e1}}\left(\Theta\right)\hat{e}_{\mathrm{1}}$. Since RATs are most efficient for $\Theta=0^{\circ}$ the data presented in this paper represents merely an upper limits for the modeling of the spin-up process of ballistic aggregates. However, focusing on the special case of $\Theta=0^{\circ}$ does not impact our comparison with previous attempts of modeling the efficiency of RATs. Simulation results are averaged over 21 equidistant values of the rotation angle of the grain around $\hat{a}_{1}$. The results are computed for 50 wavelengths between $0.05\,\mu\mathrm{m}$ and $150\,\mu\mathrm{m}$ within equidistant bins in log-space. Each combination of material, monomer-density, size, and shape (see Sect. 2) is simulated with five different resolutions i.e. the separation of two nearest dipoles. Those are combined using an average weighted by dipole resolution. In addition to the RAT efficiency factor ${Q}_{\mathrm{\Gamma}}$, DDSCAT also provides the efficiency of absorption ${Q}_{\mathrm{abs}}$ required to evaluate the photon-drag-timescale $\tau_{\mathrm{IR}}$ and the dust temperature $T_{\mathrm{dust}}$. The gas-torque-efficiency $Q_{\mathrm{gas}}$ needed to evaluate the gas-drag-timescale $\tau_{\mathrm{gas}}$ is calculated with a Monte Carlo (MC) based gas-dust collision approach as outlined in Reissl et al. (2023). The corresponding gas parameters [^4] We note that the gas number density $n_{\mathrm{gas}}$ is lower than in a typical AGN torus. However, we can position the dust grain outside the torus, selecting a position in which $n_{\mathrm{gas}}$ is similar to the ISM, e.g. at higher inclinations from the plane of the torus. are listed in Tab. 1.
Table 3: Range and mean of individual grain dust temperatures
source $T_{\mathrm{min}}$ [K] $T_{\mathrm{mean}}$ [K] $T_{\mathrm{max}}$ [K]
CA
AGN 368 410 471
S2 160 172 189
S1 114 121 130
ISRF 72 79 88
SC
AGN 119 136 161
S2 37 41 48
S1 22 24 26
ISRF 15 16 18
@@ 6.1 Drag timescales
Resulting timescales for the gas-drag and the photon-drag, respectively, taking different gas densities and dust temperatures are shown in Fig. 3. The gas-drag-timescale $\tau_{\mathrm{gas}}$ is presented in the left panel as an average of SC and CA materials. We emphasize that the gas-drag-efficiency $Q_{\mathrm{gas}}$ is only dependent on grain shape in our modeling and is independent of material. We see that denser gas equals a smaller $\tau_{\mathrm{gas}}$ in agreement with the literature (see e.g. Draine & Weingartner 1997 and Reissl et al. 2023). The dependence on monomer-density is marginal, with the BA grains having a stronger gas-drag than the BAM1 and BAM2 grains. This is due to the dependence of $Q_{\mathrm{gas}}$ on the surface area.
In detail, the BA grains present a larger surface area for interaction with gas particles. Following Eq.5, the expected size dependence of $\tau_{\mathrm{gas}}$ would be linear ($I_{\mathrm{max}}\propto a_{\mathrm{eff}}^{5}$). However it is almost constant above $100\,\mathrm{nm}$, which is due to $Q_{\mathrm{gas}}$ increasing linearly with size. In principle the gas drag efficiency $Q_{\mathrm{gas}}$ should only depend on the shape of the grain and not the size. That $Q_{\mathrm{gas}}$ nonetheless increases with size may be explained by $a_{\mathrm{eff}}$ not being well suited to describe grain size in this situation. The calculation of $Q_{\mathrm{gas}}$ uses the simulated gas-drag-torque, which depends on the angular momentum lost by the grain when a gas particle leaves the surface. This in turn scales linearly with the distance from the axis of rotation of the grain from where the gas particle is ejected. For porous grains this distance is often larger than $a_{\mathrm{eff}}$. The torque simulated in this way contains the information about the true shape of the grain, however, in the formula for $Q_{\mathrm{gas}}$ the size $a_{\mathrm{eff}}$ is used, which creates the size dependence of $Q_{\mathrm{gas}}$.
The photon-drag-timescale is shown in the middle for CA grains and on the right for SC grains in Fig. 3. As the efficiency of absorption $Q_{\mathrm{abs}}$ is overall larger for SC, more photons are absorbed and emitted causing a stronger photon-drag. The number of emitted photons also depends on dust temperature, with hotter dust experiencing stronger drag (see also Tab. 3). The dependence on monomer-density is marginal, and is related to the fact that the moment of inertia is larger in the case of BA grains than for the more compact BAM2 grains given the same effective radius (see Appendix A in Reissl et al. 2023).
@@
Figure 6: The same as Fig. 5 but for the average torque efficiency $Q_{\Gamma}$ over $\lambda/a_{\mathrm{eff}}$ in comparison with the conical power-law parametrization (red dotted line) as suggested in LH07.
Figure 7: Two-power-law parametrization shown as a red line for CA (left) and SC (right) BA-grains in comparison with the full set of the numerical DDSCAT simulations (black dots) for grains of size $a_{\mathrm{eff}}=200\ \mathrm{nm}$ where the absolute value of the torque efficiency $|Q_{\Gamma}|$ are plotted as a function of $\lambda/a_{\mathrm{eff}}$. The constant part $c$ is fitted for each grain size up to a value of $\lambda/a_{\mathrm{eff}}=3$, the steeper power-law is fitted as a linear function to the logarithm of the data for all sizes combined in the interval $103<\lambda/a_{\mathrm{eff}}<$ 10, same for the shallow power-law in the interval $\lambda/a_{\mathrm{eff}}>20$. Finally, the two intersection points $a$ and $b$ are calculated and used to create a continuous parametrization for $|Q_{\Gamma}|$ (see Eq. 13).